At a car rental agency .50 of the cars are returned on time a sample of 13 car rentals is studied what is the probability that more than 3 of them are returned on time?
Answer #1

Here, n = 13, p = 0.5, (1 - p) = 0.5 and x = 3
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X <= 3).
P(X <= 3) = (13C0 * 0.5^0 * 0.5^13) + (13C1 * 0.5^1 * 0.5^12) + (13C2 * 0.5^2 * 0.5^11) + (13C3 * 0.5^3 * 0.5^10)
P(X <= 3) = 0.0001 + 0.0016 + 0.0095 + 0.0349
P(X <= 3) = 0.0461

P(X > 3) = 1 - 0.0461 = 0.9539